# Learning Program Behavioral Models from Synthesized Input-Output Pairs

### v1

@@ 6. Related Work
Neural machine translation techniques have also shown promising results in low-level code decompilation tasks. Such neural decompilers are either implemented as sequence-to-sequence (Katz et al., 2018; Liang et al., 2021) sequence-to-tree (Fu et al., 2019; Cao et al., 2022) models. While existing works focus on converting the program’s binary representation to high-level program code, our framework allows users to reverse the direction of computation (predict the program input that was used to produce a given output) by training a behavioral model of the inverse program. This result is achieved by simply swapping training data pairs, i.e., using program output as model input and program input as model output. Even though such an inverse program may not exist, our approach still allows framework users to estimate the result of an inverse computation. Meanwhile, the approaches proposed by Katz et al. in the TRAFIX framework (Katz et al., 2019) and Fu et al., in the N-Bref framework (Fu et al., 2021), suggest training a neural decompiler from generated data. We think that these works are the closest to ours. In contrast, our framework supports behavior modeling of arbitrary programs, learning both direct and inverse models of the selected PUT, and supports further generalization of the learning process from mapping finite input-output pairs to extracting common input properties and constraints.
Data and Model Poisoning.:
@@
Language models trained from crowd-sourced data suffer from data poisoning problems. In their research, Pearce et al. (Pearce et al., 2022) have studied the impact of training dataset poisoning on the quality of language model predictions. Code-generation models still tend to generate buggy or malicious code snippets because they have been trained on very similar corpora of publicly available documents. Schuster et al. (Schuster et al., 2021) have demonstrated two types of attacks on generative models: data poisoning and model poisoning. In the first scenario, only a few specially crafted samples, which are added to the training corpus, are sufficient to enforce code generation models to produce malicious code snippets. Alternatively, such models can be attacked by manipulating the training parameters (weights) or fine-tuning them on a few malicious samples. Both attack strategies will result in a deviation of generated sequences, and existing defense measures against dataset poisoning attacks cannot mitigate the problem completely (Wan et al., 2022; He and Vechev, 2023). Due to the possibility of manipulating model behavior, the model outcome cannot be trusted, and their application in mission-critical systems is still questionable. In contrast to other approaches that rely on crowd-sourced training data, our framework allows full control over the data collection and model training processes through the specification of input synthesis rules, testing of generated samples with the PUT, and filtration or sanitization of faulty, undesired, and dangerous samples.
Domain-specific Adaptations.:
@@
Due to the reason that such domain-specific models were fine-tuned only on the program source codes, they cannot capture the program behavior patterns without further fine-tuning. Such fine-tuning could be an expensive process due to the sizes of models and specific hardware requirements. Zheng et al. and their LlamaFactory citezheng2024llamafactory framework provide estimated memory requirements implied in fine-tuning open-source LLMs. For example, the small versions of commonly used LLMs, like Llama2, Llama3, and Mixtral, with 7-to-8 billion parameters require more than 60GB of memory for full-precision fine-tuning. Even though models are open-source, it is mostly impossible to fine-tune them locally without access to data-center-grade hardware. Instead, our framework can learn “lightweight” program behavior models consisting of only several million parameters and can be trained on commodity hardware.
Neural Program Induction.:
@@
We find that the neural program induction and neural program synthesis fields are also closely related to our research. While neural program induction models (Joulin and Mikolov, 2015; Devlin et al., 2017a; Neelakantan et al., 2016; Graves et al., 2014; Zaremba et al., 2016; Kurach et al., 2016) try to train a sequence-to-sequence generative model that learns data patterns from input-output examples, neural program synthesis (Vijayakumar et al., 2018; Devlin et al., 2017b; Shin et al., 2019) use neural networks for predicting algorithms from corresponding input-output samples. Existing studies cannot capture the behavior of complex algorithms and programs because they had been relying on already outdated neural network architectures; they could not generalize to arbitrary complex inputs because the corresponding models had been trained on a limited number of training samples, which, for certain studies are in human-generated; require the additional manual modeling and encoding of the algorithm or program in a domain-specific language that has limited functionality. Recent advances in sequence-to-sequence prediction algorithms (Transformer architecture and its variations) make it possible to automatically infer complex data patterns from input-output pairs. Moreover, our approach can also automatically learn the inverse patterns, which allows the revision of the computation results of certain programs at a small cost.
AI-driven Testing.:
@@
AI-driven, or specifically LLM-driven software testing is becoming more and more popular. We picked a few representative works from myriads of publications in this field to highlight the s TOGA (Dinella et al., 2022) project explores neural test oracle generation problem by fine-tuning the CodeBERT (Feng et al., 2020) model on automated inference of exceptional and assertion oracles from developer-written tests. Haluptzok et al. (Haluptzok et al., 2023) synthesize and solve Python programming puzzles using language models. The authors of DeepAnalyze (Shetty et al., 2022) learned a sequence labeling model to automatically mark memory frames in a stack trace that caused the crash of an application. The LIBRO (Kang et al., 2023) framework uses the Codex Large Language Model to generate and rank bug-reproducing test cases from bug reports. Tufano et al. (Tufano et al., 2019) train a neural machine translation model to automatically replicate in 36% of cases the changes between pull requests in code repositories like bug fixes or code refactoring. The DeepFix (Gupta et al., 2017) approach randomly mutates correct programs to introduce errors. It uses these data pairs to train the neural machine translation model for automated program repair, which succeeds in fixing buggy programs in 27% of cases. The Break-It-Fix-It (Yasunaga and Liang, 2021) approach achieves high accuracy in code repair tasks by chaining two generative neural machine translation models for synthesizing naturally looking code errors and training a fixer model from synthesized pairs. Godefroid et al., in their “Learn&Fuzz” publication (Godefroid et al., 2017), learn a generative model for PDF structures by training a recurrent neural network from a set of PDF objects. The input sequence is represented by a sequence of characters in PDF objects, while the output sequence is obtained by shifting the corresponding input sequence by one character. While every mentioned work follows its unique approach, they all require a significant amount of human effort for problem-solving. In contrast, Modelizer tries to automate the process of data collection (which is one of the most time-consuming parts of model preparation) and model training. The minimalistic set of requirements, in particular the ability to run a Python interpreter locally, simplifies the integration of the learned models into the existing projects and testing tools.
AI-generated Synthetic Data:
@@
The capabilities of LLMs revealed several approaches of their application for synthetic data generation (Li et al., 2023). They can generate human-like text or dialogues given the prompt query, augment existing datasets by producing variations of the original data, and produce structured representations of existing data like JSON or CSV. Such AI-generated synthetic data is mostly used for training classifiers or question-answering systems. For example, general-purpose LLMs like OpenAI ChatGPT or Google Bard have shown promising results (İrfan AYGÜN, 2023) in generating synthetic data in a medical text production task. However, in certain tasks, like hateful speech detection and classification, the inclusion of syntactically generated samples to the training set does not improve the performance of classifiers (Kruschwitz and Schmidhuber, 2024). Still, it is an active and promising field of research, and industry giants like Nvidia invest in the development of LLMs, like Nemotron-4 (Nvidia et al., 2024), specifically designed for synthetic data generation. While synthetic data generation with LLMs implies specific hardware requirements to run models locally or lead to a privacy compromise while processing the data with a third-party cloud provider at an additional service cost, our approach that relies on synthesizing data with grammar fuzzers only requires a system with an ability to run a Python interpreter locally.
Data Generation by Mutating Inputs:
@@
We also want to address the alternatives to grammar-based and AI-based input generation strategies. Potentially, an input set for the given PUT can be automatically constructed with the help of mutation-based fuzz-testing (Barton P. Miller et al., 1990). A mutation fuzzer, like AFL++ (Andrea Fioraldi et al., 2020), can accept one input for a PUT as an initial seed and produce infinitely many new inputs by applying pre-defined mutation strategies, like bit flips, addition, or subtraction of integers, on different parts of the given seed. Such a technique produces two types of mutants. The first type of mutants has a broken syntactical or semantic structure. These mutants will either crash the PUT if it contains a bug in its implementation, or the input parser guards will reject such input, and program execution will be terminated. The collected input-output pairs will only cover the behavior of the input parser rather than the whole program. In most cases, the second type of mutants, which would have managed to bypass the parser guards, will contain insignificant differences when compared to an initial input sample. Thus, testing the program with this kind of input will unlikely produce a diverse set of outputs. While fuzz testing is helpful for bug discovery in the implementations of input parsers, we do not think the mutation-based input generation technique is applicable to automate program behavior modeling.
Mining Input Grammars:
@@
In certain cases, input grammars can also be automatically mined from a PUT. The Mimid approach by Gopinath et al. (Rahul Gopinath et al., 2020) relies on the program instrumentation in combination with string wrappers to identify the respective control flow nodes that access the relevant portions of a program input. Kulkarni et al. presented the Arvada algorithm (Kulkarni et al., 2021) that can automatically learn a highly recursive CFG from a given program using a set of positive examples and a Boolean-valued oracle. Currently, if input specifications are not publicly available or are not recoverable from the PUT, then the user of our framework has to encode it in CFG format manually.
@@ 7. Conclusion and Future Work
With Modelizer, we provide a principled approach to turning regular programs into models. This opens up the manifold capabilities of machine learning models for several Software Engineering tasks such as mocking, testing, and debugging. The maybe most promising feature of the learned translation models is their ability to predict inputs for given outputs, not only reversing executions but also offering new opportunities for test input generation. We see Modelizer as a pioneer in project-specific learning—rather than trying to generate expensive LLMs that may or may not evolve into a one-size-fits-all, Modelizer demonstrates that it is possible to learn small models for nontrivial programs that can successfully replicate their behavior—and their reverse.
@@
One of the most promising applications of Modelizer is its use for test input generation. By treating traces or coverage as output of the PUT (and supplying an output grammar that can parse trace logs), Modelizer can learn which input features correlate with which traces or coverage features; and can then again predict which input would be required to achieve a particular trace or coverage. These inputs could then be tested automatically, and if the trace or coverage deviates, such feedback could be used to fine-tune the model (see below).
@@
Once Modelizer has extracted a model, one can also use it for monitoring the PUT, continuously comparing the program behavior against the extracted model. This could be used to detect and prevent behavior not seen during testing.
@@
In our evaluation, we have found again and again that the performance of Modelizer very much depends on the quality and diversity of the synthesized training data, and thus on the quality of the test generator. In our future work, we will experiment with smarter black-box test generators, covering grammar features and combinations thereof (Havrikov and Zeller, 2020), but also smarter grey-box fuzzing, using code coverage in the PUT as guidance (Pham et al., 2019; Aschermann et al., 2019).
@@
At this point, the models extracted by Modelizer do not consider the internal state, as, say, a server could maintain. Let us assume the PUT starts in a known initial state, and the input to the PUT is a sequence of commands. Then, having Modelizer learn the effects of command sequences should also give it the ability to detect that certain outputs depend on specific subsequences of commands (implying a specific state). Such reasoning will likely require larger models; we plan to evaluate Modelizer on more complex and stateful programs.
@@
We estimate that the model’s performance can be further improved by utilizing the feedback from the PUT, when available. In contrast to fundamental LLMs, Modelizer framework allows relatively cheap continuous re-training of the learned models. If the specifications of the target system have insignificantly changed, only affecting the system behavior, but not extending the population of the input-output vocabularies with new elements, the learned models can be fine-tuned with a few samples collected from system executions. In Section 2.4, we have represented a scenario when the PUT serves as an oracle to validate the model’s predictions. Once the model prediction does not align with execution results, the collected feedback can be passed to the model as additional training data to adjust the model’s parameters. The model has to be repeatedly re-evaluated and fine-tuned till the model’s predictions match the PUT’s execution results. We plan to study the capabilities of Modelizer and respective learned models to serve as an oracle for automatic detecting of the behavior deviation cases between various revisions and implementations of the same protocol. Also, we hope to explore the possibility of forging feedback-driven tokenizers, which will be able to automatically adjust the tokenization granularity and possibly mine the input-output grammars from black-box systems. They will rely not only on the feedback from the PUT but also leverage the feedback from the model evaluation.
1-bit models.:
Lastly, we are excited by the idea of replacing 16-bit parameter weights with a 1-bit representation (Ma et al., 2024), which significantly reduces the memory requirements, and thus allows training more cost-effective models. We plan to integrate this feature into Modelizer soon.
Modelizer and all experimental data are available at
@@
This work is funded by the European Union (ERC S3, 101093186). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@ 6.2. Reversing Computation
Neural machine translation techniques have shown promising results in low-level code decompilation tasks. Such neural decompilers are implemented as sequence-to-sequence (Katz et al., 2018; Liang et al., 2021) or sequence-to-tree (Fu et al., 2019; Cao et al., 2022) models. While existing works focus on converting a program’s binary representation to high-level code, our framework lets users reverse computation direction (predict the program input used to produce a given output) by training a model of the inverse program. This is done by simply swapping training data pairs, i.e., using program output as model input and program input as output. Even when an exact inverse program may not exist, our approach allows users to estimate the result of an inverse computation. Approaches by Katz et al. in the TRAFIX framework (Katz et al., 2019) and Fu et al. in the N-Bref framework (Fu et al., 2021) suggest training a neural decompiler from generated data, which we find closest to our work. In contrast, our framework supports behavior modeling of arbitrary programs, learning both direct and inverse models of the selected PUT. It also generalizes the learning process from mapping finite input-output pairs to extracting common input properties and constraints.
6.3. Data and Model Poisoning
@@
Language models trained on crowd-sourced data suffer from poisoning problems. Pearce et al. (Pearce et al., 2022) studied the impact of dataset poisoning on language model predictions’ quality. Code-generation models still tend to generate buggy or malicious code snippets because they are trained on very similar corpora of publicly available documents. Schuster et al. (Schuster et al., 2021) demonstrated two types of attacks on generative models: data poisoning and model poisoning. In the first scenario, only a few specially crafted samples added to the training corpus are enough to make code generation models produce malicious snippets. Alternatively, such models can be attacked by manipulating the training parameters (weights) or fine-tuning them on a few malicious samples. Both attack strategies result in a deviation of generated sequences, and existing defensive measures against dataset poisoning cannot mitigate the problem completely (Wan et al., 2022; He and Vechev, 2023). Due to possible manipulation of model behavior, the model outcome cannot be trusted, and their application in mission-critical systems remains questionable. In contrast to other approaches relying on crowd-sourced training data, our framework allows full control over data collection and model training processes by specifying input synthesis rules, testing generated samples with the PUT, and filtering or sanitizing faulty, undesired, and dangerous samples.
6.4. Pre-trained and Unsupervised Tokenizers
Instead of implementing tokenizers from scratch, one can potentially rely on existing generic pre-trained tokenizers. For example, the HuggingFace repository (Wolf et al., 2020) provides access to many community-developed natural language tokenizers through the HuggingFace API. Such tokenizers cannot properly split the program data for several reasons. First, they ignore the data-specific or computation-specific features since they cannot distinguish structure from content, which impacts the tokenization granularity. Second, some tokenizers, like Llama3 (AI@Meta, 2024) tokenizer, modify the input by prepending or appending additional characters to the produced tokens. This can break the syntactical correctness of program data and impact the quality of learned models.
To control the tokenization process while maintaining low development effort for end users of Modelizer, we provided support to train the Google SentencePiece (Kudo and Richardson, 2018) unsupervised tokenizer for the given program program data. SentencePiece learns a subword vocabulary from raw input text using a specified vocabulary size; it treats the text as a continuous stream of characters. Once the training is complete, the tokenizer uses this learned vocabulary to break new input data into token sequences. By setting the vocabulary size during training, one can control both the lengths of the resulting token sequences and the number of trainable parameters in the model. Similar to other existing unsupervised tokenizers, SentencePiece also ignores the structural representation of data at the tokenization time. To reduce the size of token vocabularies in our experiments, we used custom tokenizers adapted for the subjects we were experimenting with.
6.5. Domain-specific Adaptations
@@
Because such domain-specific models were fine-tuned only on the program source codes, they cannot capture the program behavior patterns without further fine-tuning. Our evaluation shows that LLMs pre-trained on code do not behave as drop-in replacements for existing programs. While humans can tolerate a certain degree of noise and errors in the LLM output, other programs cannot interpret cannot interpret them without failure. Additionally, their operation is significantly more costly than interactions with custom-trained models with fewer parameters and is incomparable in terms of cost-efficiency with original programs. Fine-tuning could be an expensive process as well due to the sizes of models and specific hardware requirements. Zheng et al. and their LlamaFactory citezheng2024llamafactory framework provide estimated memory requirements implied in fine-tuning open-source LLMs. For example, the small versions of commonly used LLMs, like Llama2, Llama3, and Mixtral, with 7-to-8 billion parameters, require more than 60GB of memory for full-precision fine-tuning. Even though models are open-source, it is mainly impossible to fine-tune them locally without access to data-center-grade hardware. To mitigate these memory requirements problems, one can use a post-training model weights quantization(Krishnamoorthi, 2018; Banner et al., 2019) technique that reduces the precision of model weights, thus requiring a lower number of bits to store them. Also, the low-rank adaptors(Hu et al., 2021) technique can further optimize computation and storage efficiency by introducing low-rank matrices into specific layers of LLMs like the attention mechanism to capture task-specific information by updating only these matrices during fine-tuning. However, model weight quantization can tamper with the prediction quality due to quantization errors, which may significantly impact performance in learning uncommon behavior patterns. Similarly, LoRA adapters might limit the model’s expressiveness and adaptability by constraining updates only to low-rank matrices. Instead, our framework can learn “lightweight” program behavior models consisting of only several million parameters and can be trained on commodity hardware at a low cost.
6.6. Neural Program Induction
@@
We find that the neural program induction and neural program synthesis fields are also closely related to our research. While neural program induction models (Joulin and Mikolov, 2015; Devlin et al., 2017a; Neelakantan et al., 2016; Graves et al., 2014; Zaremba et al., 2016; Kurach et al., 2016) try to train a sequence-to-sequence generative model that learns data patterns from input-output examples, neural program synthesis (Vijayakumar et al., 2018; Devlin et al., 2017b; Shin et al., 2019) use neural networks for predicting algorithms from corresponding input-output samples. Existing studies cannot capture the behavior of complex algorithms and programs because they had been relying on already outdated neural network architectures; they could not generalize to arbitrary complex inputs because the corresponding models had been trained on a limited number of training samples, which, for certain studies are in human-generated; require the additional manual modeling and encoding of the algorithm or program in a domain-specific language that has limited functionality. Recent advances in sequence-to-sequence prediction algorithms (Transformer architecture and its variations) make it possible to automatically infer complex data patterns from input-output pairs. Moreover, our approach can also automatically learn the inverse patterns, which allows for the revision of certain programs’ computation results at a low cost.
6.7. AI-driven Testing
@@
AI-driven, or specifically LLM-driven software testing, is becoming more and more popular. We picked a few representative works from myriads of publications in this field to highlight common use case scenarios. For example, the TOGA (Dinella et al., 2022) project explores neural test oracle generation problem by fine-tuning the CodeBERT (Feng et al., 2020) model on automated inference of exceptional and assertion oracles from developer-written tests. Haluptzok et al. (Haluptzok et al., 2023) synthesize and solve Python programming puzzles using language models. The authors of DeepAnalyze (Shetty et al., 2022) learned a sequence labeling model to automatically mark memory frames in a stack trace that caused the crash of an application. The LIBRO (Kang et al., 2023) framework uses the Codex Large Language Model to generate and rank bug-reproducing test cases from bug reports. Tufano et al. (Tufano et al., 2019) train a neural machine translation model to automatically replicate in 36% of cases the changes between pull requests in code repositories like bug fixes or code refactoring. The DeepFix (Gupta et al., 2017) approach randomly mutates correct programs to introduce errors. It uses these data pairs to train the neural machine translation model for automated program repair, which succeeds in fixing buggy programs in 27% of cases. The Break-It-Fix-It (Yasunaga and Liang, 2021) approach achieves high accuracy in code repair tasks by chaining two generative neural machine translation models for synthesizing naturally looking code errors and training a fixer model from synthesized pairs. Godefroid et al., in their “Learn&Fuzz” publication (Godefroid et al., 2017), learn a generative model for PDF structures by training a recurrent neural network from a set of PDF objects. The input sequence is represented by a sequence of characters in PDF objects, while the output sequence is obtained by shifting the corresponding input sequence by one character. While each work mentioned follows its unique approach, they all require a significant amount of human effort to solve problems. In contrast, Modelizer tries to automate the process of data collection (which is one of the most time-consuming parts of model preparation) and model training. The minimalistic set of requirements, in particular the ability to run a Python interpreter locally, simplifies the integration of the learned models into the existing projects and testing tools.
6.8. AI-generated Synthetic Data
@@
The capabilities of LLMs revealed several approaches of their application for synthetic data generation (Li et al., 2023). They can generate human-like text or dialogues given the prompt query, augment existing datasets by producing variations of the original data, and produce structured representations of existing data like JSON or CSV. Such AI-generated synthetic data is mostly used for training classifiers or question-answering systems. For example, general-purpose LLMs like OpenAI ChatGPT or Google Bard have shown promising results (İrfan Aygün and Kaya, 2023) in generating synthetic data in a medical text production task. However, in certain tasks, like hateful speech detection and classification, the inclusion of syntactically generated samples to the training set does not improve the performance of classifiers (Kruschwitz and Schmidhuber, 2024). Still, it is an active and promising field of research, and industry giants like Nvidia invest in the development of LLMs, like Nemotron-4 (Adler et al., 2024), specifically designed for synthetic data generation. While synthetic data generation with LLMs implies specific hardware requirements to run models locally or leads to a privacy compromise while processing the data with a third-party cloud provider at an additional service cost, our approach, which relies on synthesizing data with grammar fuzzers, only requires a system with the ability to run a Python interpreter locally.
6.9. Data Generation by Mutating Inputs
@@
We also want to address the alternatives to grammar-based and AI-based input generation strategies. Potentially, an input set for the given PUT can be automatically constructed with the help of mutation-based fuzz-testing (Miller et al., 1990). A mutation fuzzer, like AFL++ (Fioraldi et al., 2020), can accept one input for a PUT as an initial seed and produce infinitely many new inputs by applying pre-defined mutation strategies, like bit flips, addition, or subtraction of integers, on different parts of the given seed. Such a technique produces two types of mutants. The first type of mutants has a broken syntactical or semantic structure. These mutants will either crash the PUT if it contains a bug in its implementation, or the input parser guards will reject such input, terminating program execution. The collected input-output pairs will only cover the behavior of the input parser rather than the whole program. In most cases, the second type of mutants, which would have managed to bypass the parser guards, will contain insignificant differences when compared to an initial input sample. Thus, testing the program with this kind of input will unlikely produce a diverse set of outputs. While fuzz testing is helpful for bug discovery in input parser implementations, we do not think the mutation-based input generation technique is applicable to automate program behavior modeling.
6.10. Mining Input Grammars
@@
In certain cases, input grammars can also be automatically mined from a PUT. The Mimid approach by Gopinath et al. (Gopinath et al., 2020b) uses program instrumentation with string wrappers to identify control flow nodes accessing relevant portions of a program input. Kulkarni et al. presented the Arvada algorithm (Kulkarni et al., 2021) that can automatically learn a highly recursive grammar from a given program using positive examples and a Boolean-valued oracle. The GDBMiner (Eisele et al., 2025) tool by Eisele et al. can also automatically extract input grammars from binaries and executables written in any compiled language on any system architecture that can execute a GDB debugger. Finally, symbolic parsing (Bettscheider and Zeller, 2024) by Bettscheider et al. promises to statically extract accurate input grammars from parser code, without requiring any sample inputs. Currently, if input formats are not publicly available or are not recoverable from the PUT, the user of our framework must specify them as a grammar. Once an input grammar is available, Modelizer can generate inputs and learn the program behavior.
@@ 7. Conclusion and Future Work
With Modelizer, we provide a principled approach to turning regular programs into models. This opens up the manifold capabilities of machine learning models for several Software Engineering tasks such as mocking, testing, monitoring, and debugging. Program models learned from synthetic inputs effectively represent real-world program execution scenarios by learning basic patterns from program input-output relations. While the lack of knowledge of the end data feature distribution could limit accurate model training from scratch, few-shot fine-tuning on real-world examples efficiently enhances prediction accuracy. The most promising feature of the learned translation models is their ability to predict inputs for given outputs, not only reversing executions but also offering new opportunities for test input generation. We see Modelizer as a pioneer in project-specific learning—rather than trying to generate expensive LLMs that may or may not evolve into a one-size-fits-all, Modelizer demonstrates that it is possible to learn small models for nontrivial programs that can successfully replicate their behavior—and their reverse.
@@
One of the most promising applications of Modelizer is its use for test input generation. By treating traces or coverage as output of the PUT (and supplying an output grammar that can parse trace logs), Modelizer can learn which input features correlate with which traces or coverage features; and can then again predict which input would be required to achieve a particular trace or coverage. These inputs could then be tested automatically, and if the trace or coverage deviates, such feedback could be used to fine-tune the model.
@@
Once Modelizer has extracted a model from the reference program, one can also use it for monitoring revisions of PUT by continuously comparing the program behavior against the extracted model. This could be used to detect and prevent behavior not seen during testing.
@@
In our evaluation, we repeatedly found evidence that Modelizer’s performance greatly depends on the quality and diversity of synthesized training data and, thus, on the quality of the test generator. In our future work, we will experiment with smarter black-box test generators, covering grammar features and combinations thereof (Havrikov and Zeller, 2020), but also smarter grey-box fuzzing, using code coverage in the PUT as guidance (Pham et al., 2019; Aschermann et al., 2019).
@@
At this point, the models extracted by Modelizer do not consider internal state, as, say, a server might maintain. Let us assume the PUT starts in a known initial state, and the input to the PUT is a sequence of commands. Then, having Modelizer learn the effects of command sequences should enable it to detect that certain outputs depend on specific subsequences of commands (implying a specific state). Such reasoning will likely require larger models; we plan to evaluate Modelizer on more complex and stateful programs.
@@
We estimate that the model’s performance can be further improved by utilizing the feedback from the PUT when available. In contrast to fundamental LLMs, Modelizer framework allows relatively cheap continuous re-training of the learned models. If the specifications of the target system have insignificantly changed, only affecting the system behavior but not extending the population of the input-output vocabularies with new elements, the learned models can be fine-tuned with a few samples collected from system executions. In Section 2.4, we have represented a scenario when the PUT serves as an oracle to validate the model’s predictions. Once the model prediction does not align with execution results, the collected feedback can be passed to the model as additional training data to adjust the model’s parameters. The model has to be repeatedly re-evaluated and fine-tuned until its predictions match the PUT’s execution results. We plan to study the capabilities of Modelizer and respective learned models to serve as an oracle for automatic detecting of the behavior deviation cases between various revisions and implementations of the same protocol. We also hope to explore the possibility of forging feedback-driven tokenizers, which will be able to automatically adjust the tokenization granularity and possibly mine the input-output grammars from black-box systems. They will rely not only on the feedback from the PUT but also leverage the feedback from the model evaluation.
Modelizer and all experimental data, including interactive plots, are available at
@@
This work is funded by the European Union (ERC S3, 101093186). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.